# From $U(1)\times U(1)$ Symmetry Breaking to Majoron Cosmology: Insights from NANOGrav 15-year Data

### v1

@@ V.1 Relativistic degrees of freedom ($\Delta N_{\text{eff}}$)
The future CMB-S4 experiment is expected measure $N_{\text{eff}}$ with a precision of ±$\pm 0.07$ [97]. Assuming the central value to be the SM $N_{\text{eff}}$, this will result in a $2\sigma$ constraint of $\Delta\Omega_{x}h_{0}^{2}\,\leq 7.8\times 10^{-7}$ on additional sources of radiation in a BSM model, and will yield $\eta<3.56\times 10^{15}$ GeV. Hence, CMB-S4 observations will be in tension with a part of the parameter space in the GCS-only fit to the NANOGrav 15-year data.
@@ V.2 CMB anisotropies and iso-curvature perturbations
BP 3 15 1 16 $4.9\times 10^{13}$ $5.0\times 10^{13}$ $8.1\times 10^{-16}$ $1.3\times 10^{-16}$ 0.1207
BP 4 16 1 17 $2.9\times 10^{13}$ $5.0\times 10^{13}$ $1.2\times 10^{-5}$ $9.3\times 10^{-16}$ 0.1192
@@
The total DM relic can be arrived at by combining the components from Eqs. (37), (38) and (39) which would depend on various model parameters listed as $d$, $n$, $n^{\prime}$, $q$, $q^{\prime}$, $\eta$, $\eta^{\prime}$, $\eta_{eff}$, $m_{\chi}$ and $g_{\textrm{grav}}$. Among these parameters, $\eta_{eff}$ and $m_{\chi}$ can be written in terms of the other parameters following Eqs. (9) and (10). The value of $q$ has been fixed to 2 in Section II.2 and the $q^{\prime}$ can be expressed in terms of $q$, $n$ and $n^{\prime}$ utilizing conservation of the $B-L$ charge in the higher dimensional operator of the lagrangian mentioned in Eq. (6). The value of $\eta^{\prime}$ is fixed roughly at $\mathcal{O}(10^{13})$ GeV after taking the data of Table 3 into consideration when a combination of Type-$\phi^{\prime}$ LCS and Type-$\phi$ GCS-like strings fit the NANOGrav data. Furthermore, one of the parameters among $n$ and $n^{\prime}$ can be traded off with the other and $d$ using dimensional analysis of the lagrangian term in Eq. (6). That leaves us with the parameters $d$, $\eta$, $g_{\textrm{grav}}$ and $n$ (or $n^{\prime}$) that can be varied independent of each other. In what follows, we shall try to constrain this four-dimensional parameter space using the DM relic abundance using Planck data ($\Omega_{DM}h^{2}=0.120\pm 0.001$[111]).
@@
The analysis is carried out under two parallel assumptions, $n=1$ and $n^{\prime}=1$, and the respective results are displayed. The parameter space is scanned in the following ranges: $d$ in range [5, 20], $\eta$ in range [$10^{6}$, $5\times 10^{13}$] GeV and $g_{\textrm{grav}}$ in range [$10^{-d}$, 1] 444 The smallness of $g_{\textrm{grav}}$ is quite natural for one scalar field $\phi$ breaking a global $U(1)$ symmetry as argued in Section II.2. We do not explicitly show the same for the full $U(1)\times U(1)$ modified majoron model. However, as this smallness arises from the exponential suppression by the combined action ($S$) of the scalar field and gravity, it is natural to assume tiny values of $g_{\textrm{grav}}$ in the complete modified majoron model as well. for explaining the measured DM relic. The parameter space has been restricted to $m_{\chi}<10^{-15}$ eV in order to maintain the desired GW spectrum from Type-$\phi$ CSs in the nano-Hz frequency range. The lowest operator dimension that can satisfy the correct relic within the above-mentioned range is $d=16$ for the $n=1$ case, as shown in the upper panel of Fig. 7. On the other hand, the lower panel of Fig. 7 depicts that the same analysis with $n^{\prime}=1$ also yields the same number $d=16$. A few representative benchmark points that are consistent with the constraints mentioned in Section V and can also satisfy the correct DM relic abundance have been mentioned in Table 4 and further highlighted in Fig. 7. Next, we fix the $g_{\textrm{grav}}$ at $10^{-d+2}$ and plot the relic contour in the $\eta$ vs $m_{\chi}$ plane in Fig. 8. In this figure, the points that correspond to correct relic, underabundant relic, and overabundant relic are given, respectively, by black, blue and red colours. The adjacent points represent the same value of $d$.
@@
Figure 7: (Upper panel): Data points that explain the dark matter relic abundance for $\eta^{\prime}=5\times 10^{13}$ GeV and $n=1$. (Lower panel): Same as above except for $n^{\prime}=1$ instead of $n=1$.
@@
Figure 8: (Left): Data points that explain dark matter relic abundance for $\eta^{\prime}=5\times 10^{13}$ GeV, $n=1$ and $g_{\textrm{grav}}=10^{-d+2}$. The blue and red dots represent the underabundant and overabundant regions respectively while the black stars represent the measured DM relic. (Right): Same as above except for $n^{\prime}=1$ instead of $n=1$.
@@
A careful observation of Table 4 and Fig. 9 will reveal that when we fix $\eta^{\prime}=5\times 10^{13}$ GeV the correct DM relic density is obtained for $\eta\sim\eta^{\prime}$ and $m_{\chi}\sim 10^{-16}-10^{-15}$ eV. Below these values of $\eta$ and $m_{\chi}$, majoron DM is under-abundant. Our choice of $\eta^{\prime}=5\times 10^{13}$ emanated from the fit to the NANOGrav data when both Type-$\phi^{\prime}$ and Type-$\phi$ strings contribute to the GW spectrum in this model. However, to make sure that the Type-$\phi^{\prime}$ LCSs survive long enough so that the resulting GW spectrum reaches below nano-HZ, one requires $m_{\chi}<10^{-23}$ eV in the modified majoron model. Hence, clearly, for the scenario when the latest NANOGrav data is fit by both Type-$\phi^{\prime}$ and Type-$\phi$ strings, which evades all the bounds from Section V, majoron DM will be significantly under-abundant. They can only small form a fraction of the total DM.
In contrast, if one considers $10^{-23}\,\text{eV}<m_{\chi}<10^{-15}\,\text{eV}$, then the NANOGrav data still can be fit by GW emitted by only Type-$\phi$ strings (see Sections III.3 and IV for details). However, in such a scenario we require $\eta\sim\mathcal{O}(10^{15})$ GeV. We have not incorporated $\eta$ values up to $\mathcal{O}(10^{15})$ GeV in our scan in this section since this parameter space is in tension with bounds from $\Delta N_{\text{eff}}$, CMB anisotropies, and isocurvature perturbations. However, we should mention here that among the three contributions to the DM relic mentioned in Eqs. (37), (38), and (39), $\Omega_{\chi,str}h^{2}$ happens to be the dominant one for $m_{\chi}<10^{-15}$ eV, and one can neglect the other contributions. Equation (39) implies a simple relation, $\Omega_{\chi,\text{str}}h^{2}\propto\sqrt{m_{\chi}}\eta^{2}$. Using this expression, we plot the contour corresponding to the observed DM relic abundance, $\Omega_{\chi}/\Omega_{\text{DM}}=1$ (shown in pink), extending up to $\eta=10^{16}$ GeV in the $\eta$$m_{\chi}$ plane of Fig. 9.
@@ VII Summary and conclusion
Figure 9: The majoron relic contours as a fraction of the observed DM relic for various values {$10^{-6}$, $10^{-3}$, 1} are shown on the majoron mass vs $\eta$ parameter space. The light green shaded regions are disallowed from CMB anisotropy measurements for GCS (top) and LCS (bottom). The gray shaded region is disallowed by the Planck constraint on $\eta$ and $\eta^{\prime}$ respectively from pNGB emission (top) and GW emission (bottom). Further, the excluded region of the parameter space from isocurvature perturbation is shaded in blue (both top and bottom) and the orange shaded portion is disallowed by fuzzy DM constraints from Lyman-$\alpha$ (top). For more details on the constraints see the text in Section V. The $2\sigma$ credible intervals (see Table 2 and 3) of parameter space that fit the NANOGrav data for the Type-C case without (with) SMBHB are displayed in light red shaded region (black vertical stripes) in both panels.
@@
Hence, to ensure gravity-induced $d\leq 4$ operators are forbidden, a more conservative model is adopted by gauging $U(1)_{B-L}$ and introducing an additional scalar $\phi^{\prime}$ [35] on top of the model discussed above. Hence, the model extends the SM gauge group by a local $U(1)_{B-L}$ symmetry and includes an approximate global $U(1)$ symmetry. Both $\phi$ and $\phi^{\prime}$ are charged under $U(1)_{B-L}$, while only $\phi$ carries the global charge. Assigning $B-L$ charge of $q=2$ to $\phi$ enables Majorana mass generation for RHN. In contrast, the $B-L$ charge $q^{\prime}$ is left general to enable model building for various $d$-dimensional soft global symmetry-breaking operators by the Lagrangian term of Eq. (6). A few phenomenologically motivated benchmark models are shown in Table 4. We refer to this $U(1)_{B-L}\times U(1)$ model as the modified majoron model.
@@
  • With GCS-only GW source, the NANOGrav 15-year data yields a $2\sigma$ credible interval of $2.69\times 10^{15}<\eta<3.63\times 10^{15}$ GeV (see Table 2 and Fig. 1). However, this fit is disfavored relative to the SMBHB-only scenario, with a BF of $7.66\times 10^{-5}$. When GCS and SMBHB sources are combined, the GW fit improves with BF $\sim 0.65$. The inferred $2\sigma$ upper bound on the symmetry-breaking scale is $\eta<1.26\times 10^{15}$ GeV (see Table 2 and Fig. 2). The fits constrain the neutrino-sector coupling $\delta_{33}$ (cf. Eq. (3)) to the range 12.5712.5712.5712.5716.9716.9716.9716.97 in the GCS-only scenario and to $\delta_{33}<5.89$ in the GCS+SMBHB case.
@@
  • The number of DWs attached to each string is determined by the winding numbers and the structure of the symmetry-breaking operator of Eq. (6). When one of the conditions of Eq. (27) is satisfied, only one DW is attached to each string, ensuring the collapse of the string-wall system and avoiding the ‘domain wall problem’. In this paper, we consider $n=1$ and $n^{\prime}=1$ conditions only.
@@ Acknowledgements.
The authors are grateful to express their gratitude to the Department of Atomic Energy (DAE), India, for the support of Harish-Chandra Research Institute. The authors would also like to thank Ashoke Sen, Anirban Basu, Santosh Kumar Rai, Mairi Sakellariadou, Tanmay Vachaspati and Utsav Atta for useful comments and discussions.

### v2

@@ V.1 Relativistic degrees of freedom ($\Delta N_{\text{eff}}$)
The future CMB-S4 experiment is expected measure $N_{\text{eff}}$ with a precision of $\pm 0.07$ [103]. Assuming the central value to be the SM $N_{\text{eff}}$, this will result in a $2\sigma$ constraint of $\Delta\Omega_{x}h_{0}^{2}\,\leq 7.8\times 10^{-7}$ on additional sources of radiation in a BSM model, and will yield $\eta<3.56\times 10^{15}$ GeV. Hence, CMB-S4 observations will be in tension with a part of the parameter space in the GCS-only fit to the NANOGrav 15-year data.
@@ V.2 CMB anisotropies and iso-curvature perturbations
Apart from the constraints mentioned above, there exist constraints from the LIGO non-observation of stochastic GWs which can be translated to the vevs associated with this model as $\eta\lesssim 10^{17}$ GeV for GCS-scenario, and $\eta^{\prime}\lesssim 10^{16}$ GeV for type-C modified majoron scenario; both of which remain well above the region of our interest.
@@
The total DM relic can be arrived at by combining the components from Eqs. (37), (38) and (39) which would depend on various model parameters listed as $d$, $n$, $n^{\prime}$, $q$, $q^{\prime}$, $\eta$, $\eta^{\prime}$, $\eta_{eff}$, $m_{\chi}$ and $g_{\textrm{grav}}$. Among these parameters, $\eta_{eff}$ and $m_{\chi}$ can be written in terms of the other parameters following Eqs. (9) and (10). The value of $q$ has been fixed to 2 in Section II.2 and the $q^{\prime}$ can be expressed in terms of $q$, $n$ and $n^{\prime}$ utilizing conservation of the BLB-L charge in the higher dimensional operator of the lagrangian mentioned in Eq. (6). The value of $\eta^{\prime}$ is fixed roughly at $\mathcal{O}(10^{13})$ GeV after taking the data of Table 4 into consideration when a combination of Type-$\phi^{\prime}$ LCS and Type-$\phi$ GCS-like strings fit the NANOGrav data. Furthermore, one of the parameters among $n$ and $n^{\prime}$ can be traded off with the other and $d$ using dimensional analysis of the lagrangian term in Eq. (6). That leaves us with the parameters $d$, $\eta$, $g_{\textrm{grav}}$ and $n$ (or $n^{\prime}$) that can be varied independent of each other. In what follows, we shall try to constrain this four-dimensional parameter space using the DM relic abundance using Planck data ($\Omega_{DM}h^{2}=0.120\pm 0.001$[117]).
@@
The analysis is carried out under the assumption, $n=1$ and the corresponding results are displayed. The parameter space is scanned in the following ranges: $d$ in range [5, 20], $\eta$ in range $10^{6}$, $5\times 10^{13}$] GeV and $g_{\textrm{grav}}$ in range $10^{-d}$, 1]  [^5] The smallness of $g_{\textrm{grav}}$ is quite natural for one scalar field $\phi$ breaking a global $U(1)$ symmetry as argued in Section II.2. We do not explicitly show the same for the full $U(1)\times U(1)$ modified majoron model. However, as this smallness arises from the exponential suppression by the combined action ($S$) of the scalar field and gravity, it is natural to assume tiny values of $g_{\textrm{grav}}$ in the complete modified majoron model as well. for explaining the measured DM relic. The parameter space has been restricted to $m_{\chi}<10^{-15}$ eV in order to maintain the desired GW spectrum from Type-$\phi$ CSs in the nano-Hz frequency range. The lowest operator dimension that can satisfy the correct relic within the above-mentioned range is $d=16$ for the $n=1$ case, as shown in Fig. 7. A few representative benchmark points that are consistent with the constraints mentioned in Section V and can also satisfy the correct DM relic abundance have been mentioned in Table 5 and further highlighted in Fig. 7. Next, we fix the $g_{\textrm{grav}}$ at $10^{-d+2}$ and plot the relic contour in the $\eta$ vs $m_{\chi}$ plane in Fig. 8. In this figure, the points that correspond to correct relic, underabundant relic, and overabundant relic are given, respectively, by black, blue and red colours. The adjacent points represent the same value of $d$.
@@
Figure 7: Data points that explain the dark matter relic abundance for $\eta^{\prime}=5\times 10^{13}$ GeV and $n=1$.
@@
Figure 8: Data points that explain dark matter relic abundance for $\eta^{\prime}=5\times 10^{13}$ GeV, $n=1$ and $g_{\textrm{grav}}=10^{-d+2}$. The blue and red dots represent the underabundant and overabundant regions respectively while the black stars represent the measured DM relic.
@@
A careful observation of Table 5 and Fig. 9 will reveal that when we fix $\eta^{\prime}=5\times 10^{13}$ GeV the correct DM relic density is obtained for $\eta\sim\eta^{\prime}$ and $m_{\chi}\sim 10^{-16}-10^{-15}$ eV. Below these values of $\eta$ and $m_{\chi}$, majoron DM is under-abundant. Our choice of $\eta^{\prime}=5\times 10^{13}$ emanated from the fit to the NANOGrav data when both Type-$\phi^{\prime}$ and Type-$\phi$ strings contribute to the GW spectrum in this model. However, to make sure that the Type-$\phi^{\prime}$ LCSs survive long enough so that the resulting GW spectrum reaches below nano-HZ, one requires $m_{\chi}<10^{-23}$ eV in the modified majoron model. Hence, clearly, for the scenario when the latest NANOGrav data is fit by both Type-$\phi^{\prime}$ and Type-$\phi$ strings, which evades all the bounds from Section V, majoron DM will be significantly under-abundant. They can only form a small fraction of the total DM.
[…]
@@ VII Summary and conclusion
Figure 9: The majoron relic contours as a fraction of the observed DM relic for various values $10^{-6}$, $10^{-3}$, 1} are shown on the majoron mass vs $\eta$ parameter space. The light green shaded regions are disallowed from CMB anisotropy measurements for GCS (top) and LCS (bottom). The gray shaded region is disallowed by the Planck constraint on $\eta$ and $\eta^{\prime}$ respectively from pNGB emission (top) and GW emission (bottom). Further, the excluded region of the parameter space from isocurvature perturbation is shaded in blue (both top and bottom) and the orange shaded portion is disallowed by fuzzy DM constraints from Lyman-$\alpha$ (top). For more details on the constraints see the text in Section V. The $2\sigma$ credible intervals (see Table 3 and 4) of parameter space that fit the NANOGrav data for the Type-C case without (with) SMBHB are displayed in light red shaded region (black vertical stripes) in both panels.
@@
Hence, to ensure gravity-induced $d\leq 4$ operators are forbidden, a more conservative model is adopted by gauging $U(1)_{B-L}$ and introducing an additional scalar $\phi^{\prime}$ [39] on top of the model discussed above. Hence, the model extends the SM gauge group by a local $U(1)_{B-L}$ symmetry and includes an approximate global $U(1)$ symmetry. Both $\phi$ and $\phi^{\prime}$ are charged under $U(1)_{B-L}$, while only $\phi$ carries the global charge. Assigning BLB-L charge of $q=2$ to $\phi$ enables Majorana mass generation for RHN. In contrast, the BLB-L charge $q^{\prime}$ is left general to enable model building for various $d$-dimensional soft global symmetry-breaking operators by the Lagrangian term of Eq. (6). A few phenomenologically motivated benchmark models are shown in Table 5. We refer to this $U(1)_{B-L}\times U(1)$ model as the modified majoron model.
@@
  • With GCS-only GW source, the NANOGrav 15-year data yields a $2\sigma$ credible interval of $2.69\times 10^{15}<\eta<3.63\times 10^{15}$ GeV (see Table 3 and Fig. 1). However, this fit is disfavored relative to the SMBHB-only scenario, with a BF of $7.66\times 10^{-5}$. When GCS and SMBHB sources are combined, the GW fit improves with BF $\sim 0.65$. The inferred $2\sigma$ upper bound on the symmetry-breaking scale is $\eta<1.26\times 10^{15}$ GeV (see Table 3 and Fig. 2). The fits constrain the neutrino-sector coupling $\delta_{33}$ (cf. Eq. (3)) to the range 12.5712.5716.9716.97 in the GCS-only scenario and to $\delta_{33}<5.89$ in the GCS+SMBHB case.
@@
  • The number of DWs attached to each string is determined by the winding numbers and the structure of the symmetry-breaking operator of Eq. (6). When one of the conditions of Eq. (27) is satisfied, only one DW is attached to each string, ensuring the collapse of the string-wall system and avoiding the ‘domain wall problem’. In this paper, we only consider the $n=1$ condition.
@@ Acknowledgements.
The authors are grateful to express their gratitude to the Department of Atomic Energy (DAE), India, for the support of Harish-Chandra Research Institute. The authors would also like to thank Ashoke Sen, Anirban Basu, Santosh Kumar Rai, Mairi Sakellariadou, Tanmay Vachaspati, K.S. Babu and Utsav Atta for useful comments and discussions.